What is transport phenomena ? Obtain the
expression for viscosity of gas by assuming that the number of particles crossing the plane from either side per unit area per second is 1/4 nv. Discuss its temperature and pressure dependence.
1
Expert's answer
2019-05-09T11:24:20-0400

All irreversible processes of statistical nature stemming from therandom continuous motion of molecules are called as transport phenomena. The heat conduction (energy transfer), fluid flow (momentum transfer), molecular diffusion (mass transfer), radiation and electric charge transfer in semiconductors are the examples of transport processes.


The coefficient of viscosity is given by formula


"ƞ=\frac {1} {3} \times {ulρ}"

where u is the average speed of thermal movement of gas molecules, l is the free path of molecule, ρ is the gas density


The viscosity increases with temperature because the average speed of thermal movement of gas molecules u depends on the temperature T:


"u= \sqrt{T}"

The viscosity don’t depend on the gas pressure because the average speed of thermal movement of gas molecules u don’t depend on the gas pressure.